IPP Progress Report for the week 2010.02.08 - 2010.02.12

(Up to IPP Progress Reports)

Eugene Magnier

Heather Flewelling

  • various amounts of reverting/kicking of various labels when they needed it.
  • queued up SVS test set for mops.
  • burntooled data from 01-01-2010 through about 02-08-2010, version 14 of burntool.
  • gave talk on demo month + how the ifa can access the data at astrocoffee
  • addstar automation checked in, but untested.
  • more magic test with newest version of magic (magictest%20100210% label)
  • made 104 histograms, sliced by survey, band, and survey and band, for various iterations of magic tests.

Bill Giebink

  • worked with Sturdevant AC technician to try to get PS2 AC working
  • looked more at PS2 dome/AC control code
  • collected, reduced and posted DIMM data
  • operated RoboDIMM Thursday and Friday night

Roy Henderson

  • some time looking into psLib optimsation: made timings, wrote new read function, played with struct alignment
  • max/min of object IDs now determined by ippToPsps and passed to DXLayer in the batch manifest. Other metedata may follow.
  • made improvements to DXLayer fast-FITS-read function
  • access to PSPS maui cluster and JHU cluster/SVN for DXLayer testing, Started to look at DXLayer code, as I will be testing it this week coming
  • began documentation on the IPP wiki, including an outline of the IPP to PSPS system and outstanding issues

Paul Price

  • Sent Run 5 MD detections to Dave Monet via publish system
  • Fix ppSub to put FWHM in diff inverse headers
  • Fix psastro to put zpt_obs in statistics file
  • Pattern I/O implemented and tested
  • Fixed pattern row correction list for XY15
  • Fixed SEGVs in pswarp, ppStack
  • Implemented mechanism for destreaking of background-preserved (i.e., no pattern or background subtraction) images
  • Fixed stack target PSF calculation to allow slightly bad PSFs (PAR_7 > 5e-2), though the target PSF itself is constrained to be good (PAR_7 >= 0).
  • Scheduled monthly videocons with UK partners: Wed 8am.
  • Fixed stack photometry: PSF parameters were being allowed to range without bounds
  • Running background-preserved M31 data

Bill Sweeney

  • moved data store and postage stamp server installation off of ipp049 to ippc17. Verified operation
  • worked with gavin to configure mysql instance on ippc17 correctly
  • made some changes to dsreg to improve error handling
  • fixed some bugs in the ippTools related to distribution
  • added feature to data store to hide filesets without actually deleting them
  • cleaned up old distribution bundles
  • modified addtool to insure that a given camRun's smf file is queued for a given dvodb no more than once
  • set up dvo catdir configuration
  • queued addRun's by hand and monitored construction of DVO db ThreePi?. All runs from the first 3 days of demo month have been added. Copied db to rsync server on ipp0222 and announced to ipp-users. It has at least been downloaded to MPIA.
  • started working on optimized update (Allowing update processing on a subset of a run's components. This is probably doable but will probably affect a hundred or so source files.

Chris Waters

  • Diffraction spike masking: updated diffraction spike mask rules to fix angular shift between old masks and the spikes, and updated the spike length model.
  • Nightly Science automation: expanded to cover more than just MD fields. New config file to define which surveys are processed via the automation code, and to remove values hard coded into the script.
  • Stack-stack diffs: Updated difftool to support stack stack diffs. There still seem to be a number of bugs to fix.
  • Difference automation: added warp-stack and stack-stack diff modes to the survey.pro pantasks module.